# Multiple Invertible and Partial-Equivariant Function for Latent Vector Transformation to Enhance Disentanglement in VAEs

### v1

@@ 3.1.3 Equivariance Property with Symmetric Matrix Exponential
To enhance the equivariance of L2L transformation, we set ${\bm{M}}$ of $\psi(\cdot)$ to a symmetric matrix. We show that 1) a group with the constraint guarantees equivariance of $\psi(\cdot)$ over the specific group, 2) $\psi(\cdot)$ being equivariant over subset of symmetries between the input space and transformed latent vector space, and 3) the constraint increases the probability of $\psi(\cdot)$ to be in the group (equal to be equivariant over the subset of symmetries).
@@
We particularly call the transformations as symmetries Higgins et al. (2022) to distinguish them from IPE- and I2L-transformations. For the generality of our method, we consider an arbitrary VAE model that has no restriction on creating intersections to any set as Figure 2.
In the next, we show that matrix exponential with symmetric matrix partially preserves encoder equivariance condition better than other matrices.
@@
Proof. The group $G_{S}$ is closed to matrix multiplication, and its element is always a symmetric matrix by definition. Then, any two elements in $G_{S}$ are commutative because if matrix multiplication of two symmetric matrices is symmetric then both are commutative. As a result, $\psi_{G_{S}}(\cdot)$ and group elements of $G_{S}$ are commutative ($G_{S}$ is an abelian group). Because of the commutativity, $\psi_{G_{S}}(g_{s}\circ{\bm{z}})=\mathbf{e}^{{\bm{S}}}g_{s}{\bm{z}}=g_{s}\mathbf{e}^{{\bm{S}}}{\bm{z}}=g_{s}\circ\psi_{G_{S}}({\bm{z}})$ for $g_{s}\in G_{S}$ if the group action $\circ$ is set to matrix multiplication, where $\psi_{G_{S}}\in G_{S}$. This equation satisfies the general definition of an equivariant function that a function $f(\cdot)$ is equivariant if $f(g\circ{\bm{z}})=g\circ f({\bm{z}})$ for all $g$ in a group $G$ by matching $f$, $g$, and $G$ to $\psi_{G_{S}}$, $g_{s}$, and $G_{S}$, respectively. $\blacksquare$
@@
If $q_{\phi}$ is equivariant over defined on group of symmetries $G_{I}^{a}$ and $G_{L}^{a}$, then $\psi_{G_{S}}(q_{\phi}(\cdot))$ is equivariant to symmetries in $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$ by the equivariance of $q_{\phi}$.
Proof. The function $\psi_{G_{S}}(\cdot)$ is an equivariant function over group elements in $G_{S}\cap G_{L}$ by Proposition 3.1. Then, the composite function, $\psi_{G_{S}}(\cdot)$ and $q_{\phi}$, is an equivariant function of $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$. Let $g_{L}^{a}$ be a group element in $G_{S}\cap G_{L}$, and $g_{I}^{a}$ is a group element in $G_{I}$ corresponding to $G_{S}\cap G_{L}$. More details are in Appendix C.1.
$\blacksquare$
Figure 3: Equivariant map: $\mathcal{X}$, $\mathcal{Z}$, and $\hat{\mathcal{Z}}$ are input space, latent vector space, and transformed latent vector space by L2L transfomration function $\psi(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. respectively. ${\bm{x}}\in\mathcal{X}$, ${\bm{z}}\in\mathcal{Z}$, and $\hat{{\bm{z}}}\in\hat{\mathcal{Z}}$.
Then,
@@
Proof. All $\mathbf{e}^{{\bm{S}}}\in E_{S}$ are in $E_{M}$ since $Sym_{n}(\mathbb{R})\subset M_{n}(\mathbb{R})$. However, $E_{M}\not\subset E_{S}$ because $\mathbf{e}^{{\bm{S}}}$ is always symmetric, but $\mathbf{e}^{{\bm{M}}}$ can be an asymmetric matrix. Therefore $E_{M}\not\subset E_{S}$. Therefore, the probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{S})}$ is greater than $Pr(\psi_{E_{M}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{M})}$. In the same way, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(M_{n}(\mathbb{R}))}$ because $E_{M}\subset M_{n}(\mathbb{R})$ and non-invertible functions are only in $M_{n}(\mathbb{R})$. $\blacksquare$
Therefore, $\psi_{E_{S}}$ clearly increases the probability of preserving a certain type of equivariance compared to unrestricted $\psi$ functions.
The conditional probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})$, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})$, and $Pr(\psi_{M}(\cdot)\in G_{S})$ is changed by the distribution of the observation of $\psi(\cdot)$, which depends on the model parameters. However, the inequality $Pr(\psi_{E_{S}}(\cdot)\in G_{S})>Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})$ is not changed regardless of the distribution of observation of $\psi(\cdot)$. We empirically validate the impact of equivariance with the uncertain $P(\cdot)$ to disentanglement in Section D.3.
3.1.4 Relation Between $\psi(\cdot)$ and Disentanglement
In addition to invertible and partial-equivariant properties, our IPE-transformation also guarantees zero Hessian matrix, which enhances disentanglement without any additional loss of Peebles et al. (2020). Hessian matrix of the transformation $\nabla_{\bm{z}}^{2}\psi({\bm{z}})=\nabla_{\bm{z}}(\nabla_{\bm{z}}\mathbf{e}^{{\bm{M}}}{\bm{z}})=0$ because of the irrelevance of ${\bm{M}}$ to ${\bm{z}}$. By this property, $\psi(\cdot)$ leads that independently factorizes each dimension Peebles et al. (2020), and it injects group theory based inductive bias simultaneously. This is because the group decomposition of ${\bm{z}}$ space $G=G_{1}\times G_{2}\times\cdots\times G_{k}$ corresponds to group decomposition of the transformed latent vector $\hat{{\bm{z}}}$ space $G^{\prime}=G_{1}^{\prime}\times G_{2}^{\prime}\times\cdots\times G_{k}^{\prime}$ such that each $G_{i}^{\prime}$ is fixed by the action of all the $G_{j}$ for $j\neq i$ Yang et al. (2022); Higgins et al. (2018b). This correspondence of decomposition is expected to transfer the independence between dimensions of ${\bm{z}}$ to the space of $\hat{{\bm{z}}}$ Higgins et al. (2018a).
@@ 3.2 Exponential Family Conversion for Unknown Prior
In VAE frameworks, the Gaussian normal distribution is applied as a prior. However, a prior from data is usually unknown and may not follow the Gaussian distribution Miao et al. (2022). As a solution, we present a training procedure for VAEs to build an exponential family distribution from a latent variable of an arbitrary distribution. Then, we introduce training losses obtained from the unit IPE-transformation function and EF-conversion.
3.2.1 Elements of Exponential Family Distribution Settings
@@
First, we set sufficient statistics $T(\cdot)$, log-normalizer $A(\cdot)$, and carrier or base measure $B(\cdot)$ are deterministic functions by maximizing conjugate prior for parameter $\mathcal{\xi}$. To determine the natural parameter of posterior and prior $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$, and $\hat{\bm{\epsilon}}_{m}$, we use a natural parameter generator (NPG) designed by multi-layer perceptron Charpentier et al. (2022a). As introduced in Bishop (2006); Charpentier et al. (2022a), we assume exponential family always admits a conjugate prior:
where $B^{\prime}(\cdot)$ is a normalize coefficient and $\nu$ is evidence, and it is expressed by prior natural parameter $\mathcal{\xi}$. However, generated natural parameter $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$ is not guaranteed as the appropriate parameter of the exponential family corresponds to conjugate prior. To satisfy this condition, we assume observation is a set of independent identically distributed, then Eq. 13 is modified: $p(\mathbf{X}|\bm{\theta})=\prod_{n=1}^{N}h(\mathbf{x}_{n})\text{exp}(\bm{\theta}^{\intercal}\sum_{n=1}^{N}T(\mathbf{x}_{n})-A(\bm{\theta}))$ Bishop (2006), where observation $\mathbf{X}=\{\mathbf{x}_{1},\cdots\mathbf{x}_{N}\}$. In the next, we multiply the modified formation by the prior Eq. 1 to obtain the posterior distribution Bishop (2006) as Eq. 2.
3.2.2 Distribution Approximation As an Exponential Family
The procedure represents a posterior distribution in the exponential family by adopting the following form:
$p(\bm{\theta}|\mathbf{X},\mathcal{\xi},\bm{\nu})\propto\text{exp}(\bm{\theta}^{\intercal}(\sum_{n=1}^{N}T(\mathbf{x}_{n})+\bm{\nu}\mathcal{\xi})-A(\bm{\theta})),$ (2)
@@ 3.2.2 Distribution Approximation As an Exponential Family
where sufficient statistics $T(\cdot)$ and log-normalizer, $A(\cdot)$ are known functions, samples $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\}$ from distribution, and natural parameter of posterior $\bm{\theta}$ and of prior $\mathcal{\xi}$ Bishop (2006). The functions $T(\cdot)$, and $A(\cdot)$ are deterministic functions to maximize posterior distribution. The evidence is implemented as learnable parameters $\bm{\nu}\in\mathbb{R}^{n\times n}$. The natural parameter is generated by a multi-layer perceptron as Charpentier et al. (2022a). This general form approximating an exponential family distribution with learnable parameters can extend VAEs to use a wider distribution for latent variables by simply matching $\mathbf{X}$ to generated latent variables. After IPE-transformation, we can apply the form by using the $\hat{{\bm{z}}}_{m}$, $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$, and $\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}$ for $\mathbf{X}$, $\bm{\theta}$, and $\mathcal{\xi}$, respectively.
3.2.3 EF Similarity Loss
We added a loss to converge the unrestricted distributions of $\hat{{\bm{z}}}$ to the power density function of the exponential family by constraining the posterior maximization as:
$\begin{split}\text{maximize}&~{}\log p(\bm{\theta}_{\hat{{\bm{z}}}_{m}}|\hat{{\bm{z}}}_{m},\bm{\theta}_{\hat{\bm{\epsilon}}_{m}},\bm{\nu}_{m})~{}~{}~{}\text{s.t.}~{}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\geq 0\end{split}$ (3)
@@ 3.2.3 EF Similarity Loss
$\begin{split}\Rightarrow\mathcal{L}_{s}(\hat{{\bm{z}}}_{m},\hat{\bm{\epsilon}}_{m})&=\log p(\bm{\theta}_{\hat{{\bm{z}}}_{m}}|\hat{{\bm{z}}}_{m},\bm{\theta}_{\hat{\bm{\epsilon}}_{m}},\bm{\nu}_{m})+\bm{\lambda}_{m}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{\mathbf{x}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\end{split}$ (4)
$\begin{split}\Rightarrow\mathcal{L}_{el}\coloneqq||\nabla_{\hat{{\bm{z}}}_{m},\hat{\bm{\epsilon}}_{m},\bm{\lambda}_{m}}\mathcal{L}_{s}||_{2}^{2}=0.\end{split}$ (5)
@@
3.2.4 KL Divergence for Evidence of Lower Bound
The KL divergence of Gaussian distribution Kingma & Welling (2013) is computed using mean and variance, which are the parameters of a Gaussian distribution. To introduce a loss as the KL divergence of Gaussian distribution, we compute KL divergence of the exponential family in Eq. 2 using the learnable parameter $T(\cdot)$ and $A(\cdot)$ with given natural parameter $\bm{\theta}_{\hat{{\bm{z}}}}$ and $\bm{\theta}_{\hat{\bm{\epsilon}}}$, expressed as:
$\begin{split}\mathcal{L}_{kl}&\coloneqq D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\\ &=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{{\bm{z}}}}}A(\bm{\theta}_{\hat{{\bm{z}}}})-\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{\bm{\epsilon}}}}A(\bm{\theta}_{\hat{\bm{\epsilon}}}).\end{split}$ (6)
3.2.5 KL Divergence Calibration Loss
To reduce the error between the approximation and true matrix for the matrix exponential Bader et al. (2019), we add a loss to minimize the difference of their KL divergence measured by mean squared error (MSE) as:
$\begin{split}\mathcal{L}_{cali}=\text{MSE}(&D_{\mathrm{KL}}(q_{\phi}({\bm{z}}|{\bm{x}})||p_{\theta}({\bm{z}})),D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))),\end{split}$ (7)
@@ 3.2.5 KL Divergence Calibration Loss
3.2.6 Implicit Semantic Mask
where $\mathcal{W}$ is the weight of log-normalizer, $\lambda$ is a hyper-parameter, $\mathbf{\mu}_{|\mathcal{W}_{ij}|}$, and $\mathbf{\sigma}_{|\mathcal{W}_{ij}|}$ are the mean, and standard deviation of weight respectively. Previous work Yang et al. (2020) utilizes a semantic mask in input space directly, but we inject the semantic mask implicitly on the latent space.
3.3 Integration for Multiple IPE-Transformation and EF-Conversion
$\begin{split}\mathcal{L}_{rec}&\coloneqq\frac{1}{k}\sum_{i=1}^{k}\Bigg{[}\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg{]}=\frac{1}{k}\sum_{i=1}^{k}E_{q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}})),\end{split}$ (9)
where $\hat{{\bm{z}}}_{i}=\psi_{i}({\bm{z}})$. Therefore, we define ELBO as:
$\begin{split}\mathcal{L}^{\prime}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})&=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi,\psi_{i}}({\bm{z}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}}))-\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})||p_{\psi_{i}}({\bm{z}})).\end{split}$ (10)
However, following Eq. 10, k samples are generated, and each sample is disentangled for different factors. We implement the output as the average of the sum of the k samples to obtain a single sample with a superposition effect from k samples. Moreover, the KL divergence term in Eq. 10 represents that increasing number of MIPE-transformation is equal to increasing $\beta$ hyper-parameter in $\beta$-VAE Higgins et al. (2017).
Table 2: Performance (mean $\pm$ std) of four metrics on dSprites, 3D Shapes, and 3D Cars. The $\alpha=1$ and $\gamma=1$ of $\beta$-TCVAE as Chen et al. (2018).
dSprites Disentanglement Metric
FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
@@ 3.3 Integration for Multiple IPE-Transformation and EF-Conversion
3D Shapes Disentanglement Metric
FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
@@
3D Cars Disentanglement Metric
FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$

### v3

@@ 4.1 Partial-Equivariant Function and Invertible L2L Transformation
We prove that an invertible and symmetric matrix preserves partial-equivariance better than other matrices for Eq. 7. We show that 1) a group with symmetric and invertible matrices guarantees equivariance of $\psi(\cdot)$ over the specific group $G_{S}$, 2) this $\psi(\cdot)$ being equivariant over subset of symmetries between the input space and transformed latent vector space (equal to be a partial-equivariant), and 3) the invertible and symmetric matrix (symmetric matrix exponential) increases the probability of $\psi(\cdot)$ to be in the group (equal to satisfy Eq. 7).
@@
For the generality of our method, we consider an arbitrary VAE model that has no restriction on creating intersections to any set as Fig. 2(b).
@@
Proof. The group $G_{S}$ is closed to matrix multiplication, and its element is always a symmetric matrix by definition. Then, any two elements in $G_{S}$ are commutative because if matrix multiplication of two symmetric matrices is symmetric then both are commutative by Lemma B.1. As a result, $\psi_{G_{S}}(\cdot)$ and group elements of $G_{S}$ are commutative ($G_{S}$ is an abelian group). Because of the commutativity, $\psi_{G_{S}}(g_{s}\circ{\bm{z}})=\mathbf{e}^{{\bm{S}}}g_{s}{\bm{z}}=g_{s}\mathbf{e}^{{\bm{S}}}{\bm{z}}=g_{s}\circ\psi_{G_{S}}({\bm{z}})$ for $g_{s}\in G_{S}$ if the group action $\circ$ is set to matrix multiplication, where $\psi_{G_{S}}\in G_{S}$. This equation satisfies the general definition of an equivariant function Eq. 1 by matching $f$, $g$, and $G$ to $\psi_{G_{S}}$, $g_{s}$, and $G_{S}$, respectively. $\blacksquare$
@@
If $q_{\phi}$ is equivariant over defined on group of symmetries $G_{I}^{J}$ and $G_{L}^{J}$, then $\psi_{G_{S}}(q_{\phi}(\cdot))$ is equivariant to symmetries in $G_{I}$ corresponding to $G_{S}\cap G_{L}$ ($G_{I}^{J}$) and $G_{T}$ corresponding to $G_{S}\cap G_{L}$ ($G_{T}^{J}$) by the equivariance of $q_{\phi}$.
Proof. The function $\psi_{G_{S}}(\cdot)$ is an equivariant function over group elements in $G_{S}\cap G_{L}$ by Proposition 4.1. Then, the composite function, $\psi_{G_{S}}(\cdot)$ and $q_{\phi}$, is an equivariant function of $G_{I}^{J}$ and $G_{T}^{J}$. Let $g_{L}^{J}\in G_{L}^{J}$, and $g_{T}^{J}\in G_{T}^{J}$. Then, $\psi_{G_{S}}(g_{L}^{J}{\bm{z}}_{1})=g_{L}^{J}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{J}\psi_{G_{S}}({\bm{z}}_{1})$. $\blacksquare$
Therefore, the invertible and symmetric matrix is a partial-equivariant function between the data space and transformed latent vector space. More details are in Proposition B.2.
Let $P(G):=\frac{\mu(G\cap K)}{\mu(K)}$ by the assumption 4 in section 4, and $Pr(\psi_{G^{\prime}}(\cdot)\in G)=P(\psi(\cdot)\in G|\psi(\cdot)\in G^{\prime})$ after VAE training.
@@
Proof. All $\mathbf{e}^{{\bm{S}}}\in E_{S}$ are in $E_{M}$ since $Sym_{n}(\mathbb{R})\subset M_{n}(\mathbb{R})$. However, $E_{M}\not\subset E_{S}$ because $\mathbf{e}^{{\bm{S}}}$ is always symmetric, but $\mathbf{e}^{{\bm{M}}}$ can be an asymmetric matrix. Therefore $E_{M}\not\subset E_{S}$. Therefore, the conditional probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{S})}$ is greater than $Pr(\psi_{E_{M}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{M})}$. In the same way, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(M_{n}(\mathbb{R}))}$ because $E_{M}\subset M_{n}(\mathbb{R})$ and non-invertible functions are only in $M_{n}(\mathbb{R})$. More details are in Proposition B.3. $\blacksquare$
Therefore, $\psi_{E_{S}}$ clearly increases the probability of Eq. 7, because the group $G_{L}$ is obtained ($|\hat{\mathcal{Z}}^{J}(\psi)|$ is fixed) by our assumption that $q_{\phi}(\cdot)$ is a partial-equivariant. It implies that symmetric and invertible properties for a matrix are necessary to preserve partial equivariance.
The probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})$, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})$, and $Pr(\psi_{M}(\cdot)\in G_{S})$ is changed by the distribution of the observation of $\psi(\cdot)$, which depends on the model parameters. However, the inequality $Pr(\psi_{E_{S}}(\cdot)\in G_{S})>Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})$ is not changed regardless of the distribution of observation of $\psi(\cdot)$. We empirically validate the impact of equivariance with the uncertain $P(\cdot)$ to disentanglement in Section D.5.
Invertible Property by Using Matrix Exponential
To guarantee the invertible property of IPE-transformation, we use a function $\psi(\cdot)=\mathbf{e}^{{\bm{M}}}\ast\cdot$ for the transformation, where $\bm{M}$} is in $n\times n$ real number matrix set $M_{n}(\mathbb{R})$ (Xiao and Liu, 2020). The operator $\ast$ is matrix multiplication, and $\mathbf{e}^{{\bm{M}}}=\sum_{k}^{\infty}\frac{{\bm{M}}^{k}}{k!}$. Our motivation is to use the benefits of injecting explicit inductive bias for disentanglement (Locatello et al., 2019; Miao et al., 2022). InteL-VAE effectively extracts hierarchical representation, which includes low-level features (affect to a specific factor) and high-level features (affect to complex factors) with an invertible transformation function (Miao et al., 2022).
@@ 4.2 Exponential Family Conversion for Unknown Prior
In VAE frameworks, the Gaussian normal distribution is applied as a prior. However, a prior from data is usually unknown and may not follow the Gaussian distribution (Miao et al., 2022). As a solution, we present a training procedure for VAEs to build an exponential family distribution from a latent variable of an arbitrary distribution. Then, we introduce training losses obtained from the unit IPE-transformation function and EF-conversion.
Elements of Exponential Family Distribution Settings
@@
As introduced in Bishop (2006); Charpentier et al. (2022), we assume exponential family always admits a conjugate prior Eq. 5. To determine the natural parameter of posterior and prior $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$, and $\hat{\bm{\epsilon}}_{m}$, we use a natural parameter generator (NPG) designed by multi-layer perceptron (Charpentier et al., 2022). However, generated natural parameter $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$ is not guaranteed as the appropriate parameter of the exponential family corresponds to conjugate prior. To satisfy this condition, we assume observation is a set of independent identically distributed, then we follow Eq. 4.
Distribution Approximation As an Exponential Family
@@
By our exponential family settings, a posterior distribution in the exponential family by adopting the following Eq. 6. The evidence is implemented as learnable parameters $\bm{\nu}\in\mathbb{R}^{n\times n}$, and natural parameter is generated by a multi-layer perceptron as Charpentier et al. (2022). This general form approximating an exponential family distribution with learnable parameters can extend VAEs to use a wider distribution for latent variables by simply matching $\mathbf{X}$ to generated latent variables. After IPE-transformation, we can apply the form by using the $\hat{{\bm{z}}}_{m}$, $\bm{\theta}_{\hat{{\bm{z}}}_{m}}$, and $\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}$ for $\mathbf{X}$, $\bm{\theta}$, and $\mathcal{\xi}$, respectively as shown in Fig. 1.
EF Similarity Loss
To maximize Eq. 6 to be an exponential family, we added a loss through a Lagrange multiplier as:
$\begin{aligned} \text{maximize}&~\log p(\bm{\theta}_{\hat{{\bm{z}}}_{m}}|\hat{{\bm{z}}}_{m},\bm{\theta}_{\hat{\bm{\epsilon}}_{m}},\bm{\nu}_{m})~~~\text{s.t.}~D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\geq 0\end{aligned}$
@@
$\begin{split}\Rightarrow\mathcal{L}_{s}(\hat{{\bm{z}}}_{m},\hat{\bm{\epsilon}}_{m})&=\log p(\bm{\theta}_{\hat{{\bm{z}}}_{m}}|\hat{{\bm{z}}}_{m},\bm{\theta}_{\hat{\bm{\epsilon}}_{m}},\bm{\nu}_{m})\\ &+\bm{\lambda}_{m}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{\mathbf{x}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\end{split}$ (9)
[…]
@@
KL Divergence for Evidence of Lower Bound
We compute the KL divergence of the exponential family in Eq. 9 using the learnable parameter $T(\cdot)$ and $A(\cdot)$ with given natural parameter $\bm{\theta}_{\hat{{\bm{z}}}}$ and $\bm{\theta}_{\hat{\bm{\epsilon}}}$, expressed as:
[…]
Further details are in the Appendix B.1.
KL Divergence Calibration Loss
Although $\mathrm{KL}$ is invariant under a common invertible map $\psi$ (Miao et al., 2022), practical approximations (e.g., matrix-exponential) can break this. We therefore add a consistency regularizer that penalizes the gap between empirical KLs before and after applying $\psi$:
$\begin{aligned} \mathcal{L}_{\text{cali}}=\text{MSE}(D_{\mathrm{KL}}(q_{\phi}({\bm{z}}|{\bm{x}})\,||\,p_{\theta}({\bm{z}})),D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})\,||\,f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}})))\end{aligned}$
@@
Table 1: Performance (mean $\pm$ std) of four metrics on dSprites, 3D Shapes, and 3D Cars.
dSprites FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
@@
3D Shapes FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
@@
3D Cars FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$